# Are AI-Generated Fixes Secure? Analyzing LLM and Agent Patches on SWE-bench

### v2

@@ II-E Metrics for Empirical Analysis
II-E1 RQ1 & RQ2: Vulnerability Prevalence and Properties
@@
II-E2 RQ3: Code, Issue, Project Characteristics Associated with LLM or agent-generated Vulnerabilities
We investigate whether the context in which the model is tasked with resolving an issue influences the security of the generated code. We analyze a range of metrics that have been shown in prior work to correlate with the riskiness of code contributions [46, 47, 48, 49], as well as introduce new metrics. Overall, these factors reflect different dimensions of task complexity and clarity. For instance, we hypothesize that issues that clearly and extensively describe the problem can be easier for LLMs to address, while tasks involving large or complex code changes may introduce more room for errors. In addition, characteristics of the project, such as size, complexity, and developer activity, may affect the likelihood of producing vulnerable code [50]. We perform this analysis for all vulnerabilities introduced both by Llama and the agentic workflows, allowing us to identify shared and distinct risk factors across different generation techniques.
@@ II-E2 RQ3: Code, Issue, Project Characteristics Associated with LLM or agent-generated Vulnerabilities
# Files added or modified: A larger numbers of files can indicate broader changes and a higher overall risk [46, 47, 48, 49].
# Unique file types modified: Modifying diverse file types may introduce different complexities to the issue resolution process [49, 46].
# Sensitive file modifications: Inspired by prior work on anomaly detecting malicious commits [49], we define a list of sensitive file types (e.g., .conf, .yml, .json) whose modification can increase the risk of vulnerabilities.
@@
Cyclomatic Complexity of LLM-Generated Code: This metric reflects the complexity of the functions introduced or rewritten by the LLM. Prior studies also analyzed the cyclomatic complexity of LLM-generated code [14, 51]. Higher complexity in generated code suggests that the model is constructing more control flows, which can be more error-prone and harder to verify for correctness and security.
@@
Issue Type and Bug Type Classification: We manually label whether the issue is a bug, a feature request, or a question, following established comprehensive taxonomies [52, 53]. Further, for issues categorized as bugs, we classify the bug type using a comprehensive, empirically grounded, and widely-adopted taxonomy[54]. The bug categories are: Configuration issues, Network issues, Database-related issues, GUI-related issues, Performance issues, Permission/deprecation issues, Security issues, Program anomaly issues, Test code-related issues, and Miscellaneous/other issues.
Information Completeness: Based on prior work emphasizing the importance of issue clarity [53], we manually annotate whether each issue includes: (1) Expected Behavior: how the software is supposed to behave, (2) Observed Behavior: how it actually behaves, and (3) Steps to Reproduce: the process needed to trigger the issue. The presence of this information can aid in understanding the bug’s context and identifying the root cause of the problem.
@@
Presence of Code Snippets: Using regular expressions, we identify code snippets in issues and use it as an indicator of technical detail and potential guidance for bug resolution.
Manual annotation of issue type, bug type, and information completeness was performed independently by two annotators, each with at least three years of programming experience. For each data point, they examine both textual and code-related data, including the GitHub issue title, description, discussion thread, etc. The annotators followed a shared set of annotation instructions (included in our replication package). They followed an iterative analysis comprising multiple sessions. Inter-rater agreement was initially measured using Cohen’s kappa for each category (all values $>$ 0.61). The annotation process proceeded through several rounds of labeling and discussion, during which disagreements were resolved. This resulted in a final strong agreement level (Cohen’s kappa values $>$ 0.9).
Project-related Characteristics: Finally, we assess project characteristics associated with the vulnerable LLM patches:
@@
# Files: The number of files in a given project can serve as a proxy for its potential complexity and size. Prior work on open-source software effort estimation has used file count as an indicator of project scale and structural complexity [55]. Further, size of a project has been shown to have direct associations with the number of potential vulnerabilities [50].
Cyclomatic Complexity and Maintainability Index of Project: We use Radon [56], a popular code analysis tool [57, 58], to compute each project’s average maintainability index and cyclomatic complexity. Cyclomatic complexity is computed as the average of complexity scores across all functions in the codebase. Higher values indicate more intricate control logic throughout the project, potentially making it harder for LLMs to reason about and modify the code securely.
@@
III Results
III-A RQ1: Security of patches generated by developers vs. standalone LLM
Vulnerability Prevalence: Table I presents the number of files with new vulnerabilities introduced by Developer and Llama 3.3 patches. The raw counts vary significantly across tools and code sources. For the full SWE-bench dataset (train+test), Llama 3.3 generated 1440 vulnerabilities, of which 185 (20.1%) passed the majority vote filter. In contrast, developer patches for the full dataset, triggered 1398 alerts (primarily driven by Bandit’s high sensitivity) but only 20 vulnerabilities (2%), identified in 17 files, were retained after majority vote filtering.
Source Semgrep CodeQL Bandit Majority Vote
Llama 3.3 (Train+Test) 323 62 1055 185
Developers (Train+Test) 9 41 1348 17
OH (Test) 447 64 1383 195
ACR (Test) 5 3 12 3
HC (Test) 19 3 37 2
TABLE I: Number of vulnerabilities detected per tool. Majority vote presents the number of files with at least one vulnerability.
When comparing individual tools, Bandit reported the highest number of vulnerabilities overall: 1383 in LLaMa-generated code and 1055 in developer code. Semgrep detected 323 vulnerabilities in LLM outputs but just 9 in developer code. CodeQL, the most conservative tool, identified 62 vulnerabilities in LLM-generated code and 41 in developer code. Overall, our results suggest that Bandit produces most false positives, consistent with prior work [59].
It is worth noting that all reported numbers are the new vulnerabilities that were injected only after the LLM/Developers modified the files or added the code to the repository. These vulnerabilities did not exist in the code before the patch generation. For instance, in an issue from the Qiskit project, Llama introduced a CWE-95 vulnerability by using eval() to interpret user-provided input strings as Python code. These strings are strictly sanitized in developers’ code. The LLM-generated code, however, converted the string into a Python lambda function without proper sanitization or sandboxing. This enabled the possibility of a code injection attack, since malicious input could be executed directly.
Vulnerability Types and Overlaps: As shown in Table II, the most common vulnerabilities in Llama-generated patches include command injection (CWE-78, 97 instances), eval injection (CWE-95, 54), insecure deserialization (CWE-502, 21), path traversal (CWE-22, 19), and weak cryptography (CWE-327, 12). While less frequent, other vulnerability types like incorrect permission assignment or error message exposure also appeared in the LLM patches. The full set of vulnerabilities are included in our replication package.
Llama-Generated Code
CWE-78 (OS Command Injection) 97
CWE-95 (Eval Injection) 54
CWE-502 (Insecure Deserialization) 21
CWE-22 (Path Traversal) 19
CWE-327 (Weak Crypto) 12
@@ III Results
CWE-732 (Permission Misconfiguration) 4
CWE-502 (Insecure Deserialization) 3
@@
CWE-377 (Insecure Temporary File) 3
CWE-703 (Exception Handling) 3
TABLE II: Most recurring CWE types found in Llama-generated and developer-written patches across the train & test sets.
Developer-written patches showed far fewer vulnerabilities overall, but some of the top categories i.e., command injection (CWE-78, 4), deserialization (CWE-502, 3), and path traversal (CWE-22, 3), are similar to the LLM results. These three common CWEs (mostly related to improper input handling and insufficient sanitization when interacting with system commands or file paths) and their presence in both sets may stem from the fact that LLMs are trained on human code, inheriting common patterns, including insecure ones.
Despite the similarities in vulnerability patterns, certain categories were far more prominent in the Llama-generated code and entirely absent from developer-written patches. Notably, eval injection (CWE-95, n=54 instances) and the use of weak cryptographic algorithms (CWE-327, n=12) ranked among the most frequent LLM vulnerabilities. These issues likely stem from the model’s tendency to translate task instructions into overly literal code, for example, using eval() to dynamically execute strings, or defaulting to insecure algorithms like MD5. Unlike developers, who often follow project standards or rely on secure external frameworks, standalone LLMs may prefer simpler but less secure implementations. This behavior reflects the model’s narrow focus on completing the task at hand with simpler code, often at the expense of secure best practices. For instance, in an issue requesting improved password hashing, Llama generated a new custom hasher that used SHA-256 instead of upgrading the existing Argon2-based implementation. While SHA-256 is straightforward to apply, it lacks protective features against brute-force attacks (CWE-327). Additionally, the LLM produced the following implementation of the check_password() function, which includes a vulnerable use of hashlib.md5():
@@
RQ1 Summary: Overall, we observed that the standalone LLM introduced up to 9× more vulnerabilities into the codebase than developers when resolving real-world issues. This highlights a significant risk associated with deploying LLMs in software development pipelines without careful security auditing. Moreover, none of the specific vulnerabilities found in Llama-generated code overlap with those in developer code (either in terms of affected files or associated issues,) indicating that LLMs can introduce distinct vulnerabilities.
III-B RQ2: Security of patches generated by the agentic APR workflows
@@
Vulnerability Prevalence: Table I also reports the number of vulnerabilities for agentic frameworks i.e., OH, ACR, and HC, when evaluated on the SWE-bench test set. Similar to RQ1, we apply the majority vote strategy to retain only high-confidence vulnerabilities.
@@
Among the agentic frameworks evaluated on the SWE-bench test set, OH produced 195 vulnerabilities (after majority vote filtering), while ACR and HC each produced 3 and 2 vulnerabilities respectively. As with Llama and developer patches, Bandit consistently reported the highest number of vulnerabilities, followed by Semgrep and then CodeQL. This pattern holds across all three frameworks. However, despite the high number of raw detections, the majority vote strategy proved highly selective: for example, OH generated over 1,383 alerts across all tools, yet only 195 vulnerabilities (approximately 15%) were retained in the final filtered set. This reinforces the importance of using a stricter intersection criterion when analyzing code security.
@@
While all frameworks introduced new vulnerabilities, OH produced a disproportionately larger number. To better understand this disparity, we closely examined the results from OH. Among other factors, OH is given full autonomy, not only to inspect and modify project files, but also to execute commands, run tests, and generate new scripts as part of its process. This broad operational scope means that it often touches more files than necessary to resolve an issue. Our analysis indicates that out of the 195 vulnerabilities initially attributed to OH, 12 were located in test files and 35 were found in files intended to reproduce the issue (e.g., reproduce.py, reproduce_error.py). Since such files are not typically intended for integration into the final production system or regular execution, we exclude them from the remainder of our analysis. Furthermore, we observe that, all the remaining 148 vulnerabilities in OH originated from just 15 issue instances. In fact, three issues alone, django ticket #27854, sphinx issue #8870, and pylint issue #4421, accounted for 124 of the 148 remaining vulnerabilities. In the most extreme case, django ticket #27854, OH generated a diff file that exceeded 500,000 lines in length. This pattern reflects a potential problem in the framework’s control mechanisms. Rather than solving the issue with minimal, targeted changes, OH can enter error-prone iterative loops that result in massive file rewrites or reimplementation of the existing core software components. These behaviors likely stem from the limitations of the agentic process in comprehending the full context of the issue and the project. As a result, OH may generate excessive or redundant code that is potentially vulnerable.
@@
In contrast, the OpenHand’s patch bypassed Django’s ORM entirely and rewrote the validation logic using raw SQL:
@@
Vulnerability Types and Overlaps: Among the agentic frameworks, OH, responsible for 148 vulnerabilities, exhibited a significantly similar vulnerability pattern to that of the standalone LLM: command injection (CWE-78, n=36), insecure deserialization (CWE-502, n=29), and eval injection (CWE-95, n=21) were all among the most common CWEs in OH patches. This alignment suggests that even when LLMs are embedded within frameworks, the core security risks and patterns persist.
@@
Moreover, we find that vulnerabilities introduced by the agentic frameworks are almost entirely different from those written by the developers. These results echo the same observations in III-A: While developers, standalone LLMs, and agentic frameworks all can add vulnerable to the codebase, the vulnerabilities produced by the LLMs or agentic frameworks are not simply repeating the developers’ mistakes.
RQ2 Summary: While the agentic frameworks were only evaluated on a smaller subset of data (test-set), all three still introduced a number of vulnerabilities. OH, in particular, produced an order of magnitude more vulnerabilities than ACR or HC. These results indicate that, despite improvements in autonomy and reasoning, current agentic workflows remain vulnerable to security issues, especially when given full control over the codebase.
III-C RQ3: What code, issue, or project characteristics are associated with the generation of vulnerable code by LLMs when resolving issues?
III-C1 Code-Level Factors
We analyze the code-level characteristics of the files associated with vulnerability-inducing LLM patches and compare them to those from all LLM-modified or generated files. Table III summarizes these comparisons for Llama-generated patches (RQ1). We note that the dataset for RQ1 is significantly larger than RQ2. As a result, due to the limited number of vulnerable examples in ACR and HC, we report statistical comparisons only for OH. However, we include the raw values for all three frameworks in Table IV.

### v3

@@ 2.5. Metrics for Empirical Analysis
2.5.1. RQ1 & RQ2: Vulnerability Prevalence and Properties
@@
2.5.2. RQ3: Code, Issue, Project Characteristics Associated with LLM or agent-generated Vulnerabilities
We investigate whether the context in which the model is tasked with resolving an issue influences the security of the generated code. We analyze a range of metrics that have been shown in prior work to correlate with the riskiness of code contributions (goyal2018identifying; alali2008s; rosen2015commit; gonzalez2021anomalicious), as well as introduce new metrics. Overall, these factors reflect different dimensions of task complexity and clarity. For instance, we hypothesize that issues that clearly and extensively describe the problem can be easier for LLMs to address, while tasks involving large or complex code changes may introduce more room for errors. In addition, characteristics of the project, such as size, complexity, and developer activity, may affect the likelihood of producing vulnerable code (gkortzis2021software). We perform this analysis for all vulnerabilities introduced both by Llama and the agentic workflows, allowing us to identify shared and distinct risk factors across different generation techniques.
@@ 2.5.2. RQ3: Code, Issue, Project Characteristics Associated with LLM or agent-generated Vulnerabilities
# Files Added or Modified: A larger number of files can indicate broader changes and a higher overall risk (goyal2018identifying; alali2008s; rosen2015commit; gonzalez2021anomalicious).
# Unique File Types Modified: Modifying diverse file types may introduce different complexities to the issue resolution process (gonzalez2021anomalicious; goyal2018identifying).
# Sensitive File Modifications: Inspired by prior work on anomaly detecting malicious commits (gonzalez2021anomalicious), we define a list of sensitive file types (e.g., .conf, .yml, .json) whose modification can increase the risk of vulnerabilities.
@@
Cyclomatic Complexity of LLM-Generated Code: This metric reflects the complexity of the functions introduced or rewritten by the LLM. Prior studies also analyzed the cyclomatic complexity of LLM-generated code (fu2023security; toth2024llms). Higher complexity in generated code suggests that the model is constructing more control flows, which can be more error-prone and harder to verify for correctness and security.
@@
Issue Type and Bug Type Classification: We manually label whether the issue is a bug, a feature request, or a question, following established comprehensive taxonomies (kallis2019ticket; song2020bee). Further, for issues categorized as bugs, we classify the bug type using a comprehensive, empirically grounded, and widely-adopted taxonomy(catolino2019not). The bug categories are: Configuration Issues, Network issues, Database-related issues, GUI-related issues, Performance issues, Permission/deprecation issues, Security issues, Program anomaly issues, Test code-related issues, and Miscellaneous/other issues.
Information Completeness: Based on prior work emphasizing the importance of issue clarity (song2020bee), we manually annotate whether each issue includes: (1) Expected Behavior: how the software is supposed to behave, (2) Observed Behavior: how it actually behaves, and (3) Steps to Reproduce: the process needed to trigger the issue. The presence of this information can aid in understanding the bug’s context and identifying the root cause of the problem.
@@
Presence of Code Snippets: Using regular expressions, we identify code snippets in issues and use them as indicator of technical detail and potential guidance for bug resolution.
Manual annotation of issue type, bug type, and information completeness was performed independently by two annotators, each with at least three years of programming experience. For each data point, they examine both textual and code-related data, including the GitHub issue title, description, discussion thread, etc. The annotators followed a shared set of annotation instructions (included in our replication package). They followed an iterative analysis comprising multiple sessions. Inter-rater agreement was initially measured using Cohen’s kappa for each category (all values >> 0.61). The annotation process proceeded through several rounds of labeling and discussion. This resulted in a final strong agreement level (Cohen’s kappa values >> 0.9). Project-related Characteristics: Finally, we assess project characteristics associated with the vulnerable LLM patches:
@@
# Files: The number of files in a given project can serve as a proxy for its potential complexity and size. Prior work on open-source software effort estimation has used file count as an indicator of project scale and structural complexity (qi2017software). Further, size of a project has been shown to have direct associations with the number of potential vulnerabilities (gkortzis2021software).
Cyclomatic Complexity and Maintainability Index of Project: We use Radon (radon), a popular code analysis tool (liawatimena2018django; hassan2024evaluating), to compute each project’s average maintainability index and cyclomatic complexity. Cyclomatic complexity is computed as the average of complexity scores across all functions in the codebase. Higher values indicate more intricate control logic throughout the project, potentially making it harder for LLMs to reason about and modify the code securely.
@@
3. Results
3.1. RQ1: Security of patches generated by Developers vs. Llama 3.3
Vulnerability Prevalence: Table 1 presents the number of files with new vulnerabilities introduced by Developers and Llama 3.3 patches. The raw counts vary significantly across tools and code sources. For the full SWE-bench dataset (train+test), Llama 3.3 generated 1440 vulnerabilities, of which 185 (12.8%) passed the majority vote filter. Upon manual inspection, 135/185 vulnerabilities remained. In contrast, developer patches for the full dataset triggered 1398 alerts, but only 17 vulnerabilities (1%) were retained after majority vote filtering. We manually verified 12/17 of these vulnerabilities to be valid security concerns.
Patch Source Split Majority Vote Manual Check
Llama 3.3 Train + Test 185 135
Developers Train + Test 17 12
Llama 3.3 Test 14 9
Developers Test 3 2
OH Test 195 44
ACR Test 3 3
HC Test 2 2
Table 1. Number of vulnerabilities detected across evaluation splits.
When comparing individual tools, Bandit reported the highest number of vulnerabilities overall: 1383 in Llama-generated code and 1055 in developer code. Semgrep detected 323 vulnerabilities in LLM outputs but just 9 in developer code. CodeQL, the most conservative tool, identified 62 vulnerabilities in LLM-generated code and 41 in developer code. Overall, our results suggest that, consistent with prior work (zhou2024comparison), Bandit produces more false positives which had to be removed in the next two steps.
It is worth noting that all reported numbers are the new vulnerabilities that were injected only after the LLM/Developers modified the files or added the code to the repository. These vulnerabilities did not exist in the code before the patches were applied. For instance, in an issue from the Qiskit project, Llama introduced a CWE-95 vulnerability by using eval() to interpret user-provided input strings as Python code. These strings are strictly sanitized in developers’ code. The LLM-generated code, however, converted the string into a Python lambda function without proper sanitization or sandboxing. This enabled the possibility of a code injection attack, since malicious input could be executed directly.
Vulnerability Types and Overlaps: Table 2 contains the most common vulnerabilities in Llama-generated and developer-written code  [^1] Each tool can assign multiple CWEs to the same finding and each vulnerability can be assigned different CWEs by different tools. Therefore the total number of CWE types is higher than the total number of vulnerabilities.. The most common vulnerabilities in Llama-generated patches include command injection (CWE-78, 79 instances), eval injection (CWE-95, 39), insecure deserialization (CWE-502, 17), and path traversal (CWE-22, 14). While less frequent, other vulnerability types like incorrect permission assignment or improper authorization also appeared in the LLM patches. The full set of vulnerabilities are included in our replication package.
Llama 3.3-Generated Code
CWE-78 (OS Command Injection) 79
CWE-95 (Eval Injection) 39
CWE-502 (Insecure Deserialization) 17
CWE-22 (Path Traversal) 14
@@ 3. Results
CWE-377 (Insecure Temporary File) 3
CWE-732 (Permission Misconfiguration) 3
@@
Table 2. Most recurring CWE types found in Llama-generated and developer-written patches across the train & test sets.
Developer-written patches showed far fewer vulnerabilities overall, but some of the top categories i.e., command injection (CWE-78, 4) and path traversal (CWE-22, 3), are similar to the LLM results. These two common CWEs (mostly related to improper input handling and insufficient sanitization when interacting with system commands or file paths) and their presence in both sets may stem from the fact that LLMs are trained on human code, inheriting common patterns, including insecure ones.
Despite the similarities in vulnerability patterns, certain categories were far more prominent in the Llama-generated code and extremely rare within developer-written patches. Notably, eval injection (CWE-95) was identified in 39 LLM patches and only one developer patch. Similarly insecure deserialization (CWE-502) was repeated in 17 LLM patches while being detected in only two developer patches. These issues likely stem from the model’s tendency to translate task instructions into overly literal code, for example, using eval() to dynamically execute strings, or defaulting to insecure algorithms like MD5. Unlike developers, who often follow project standards or rely on secure external frameworks, Llama preferred simpler but less secure implementations. This behavior reflects the model’s narrow focus on completing the task at hand with simpler code, often at the expense of secure best practices. For instance, in an issue requesting improved password hashing, Llama generated a new custom hasher that used SHA-256 instead of upgrading the existing Argon2-based implementation. While SHA-256 is straightforward to apply, it lacks protective features against brute-force attacks (CWE-327). Additionally, the LLM produced the following implementation of the check_password() function, which includes a vulnerable use of hashlib.md5():
@@
If such patches were merged without security checks e.g., CI-integrated static analysis runs or manual reviews, the consequences for a real system could be severe. In the password-hashing example, the model introduces a secondary MD5-based check that directly compares attacker-controllable input against stored hashes. In a production environment, this would expose users to significantly higher risk of credential compromise; brute-force and dictionary attacks become dramatically more feasible, and any leakage of the hash database (through a separate vulnerability) would have far more damaging consequences than under the original design. More broadly, many of the CWE categories we observe in LLM patches e.g., eval injection, insecure deserialization, and command injection, are associated with high-impact outcomes such as arbitrary code execution, privilege escalation, or unauthorized access to sensitive data. Because these vulnerabilities are introduced in the course of fixing issues, they are unlikely to be caught by functional tests alone, and can remain latent until exploited. As a result, if LLM-generated fixes are integrated without additional security checks, they may fix the immediate bug while degrading the system’s security.
RQ1 Summary: Across the SWE-bench train+test set, the standalone Llama 3.3 model introduces 11x more new vulnerabilities than developers: after majority voting and manual validation, we confirm 135 vulnerabilities in LLM-generated patches versus 12 in developer patches (Table 1). While both developers and the LLM introduce issues such as command injection (CWE-78) and path traversal (CWE-22), categories like eval injection (CWE-95) and insecure deserialization (CWE-502) are far more prevalent in LLM-generated code. Moreover, none of the confirmed LLM vulnerabilities overlap with developer vulnerabilities in the same files or issues, indicating that LLM-generated fixes can introduce distinct security weaknesses rather than merely replicating existing developer errors.
3.2. RQ2: Security of patches generated by the agentic APR workflows
@@
Vulnerability Prevalence: Table 1 also reports the number of vulnerabilities for agentic frameworks i.e., OH, ACR, and HC, when evaluated on the SWE-bench test set. Similar to RQ1, we manually inspect all instances to validate our data.
@@
The static analyses and the majority voting process resulted in 195, 3, and 2 instances marked as potential vulnerabilities by OH, ACR, and HC, respectively. As with Llama and developer patches, Bandit consistently reported the highest number of vulnerabilities, followed by Semgrep and then CodeQL. This pattern holds across all three frameworks. However, despite the high number of raw detections, the majority vote strategy proved highly selective: for example, OH generated over 1,383 alerts across all tools cumulatively, yet only 195 vulnerabilities (approximately 15%) were retained in the final filtered set. This reinforces the importance of using a stricter intersection criterion when analyzing code security. Next, led by the manual inspection process, 44 out of 195 OH vulnerabilities, and all 5 vulnerabilities from ACR and HC were confirmed as valid security threats.
@@
While all frameworks introduced new vulnerabilities, OH produced a disproportionately larger number. To better understand this disparity, we closely examined the results from OH. Among other factors, OH is given full autonomy, not only to inspect and modify project files, but also to execute commands, run tests, and generate new scripts as part of its process. This broad operational scope means that it often touches more files than necessary to resolve an issue. Our analysis indicates that out of the 195 vulnerabilities initially attributed to OH, 12 were located in test files and 35 were found in files intended to reproduce the issue (e.g., reproduce.py, reproduce_error.py). Additionally, 75 of the marked instances were parts of various packages downloaded into the repository and included as part of the patch by OH. Since such files are not typically intended for integration into the final production system or regular execution, we exclude them from the remainder of our analysis. Furthermore, we observe that, all the remaining 44 vulnerabilities in OH originated from just 11 issue instances. In fact, three issues alone, django ticket #14056, pytest issue #5787, and django ticket#14681, accounted for 35 of the 44 final vulnerabilities. In the most extreme case, django ticket #14056, OH generated a diff file that exceeded 500,000 lines in length. This pattern reflects a potential problem in the framework’s control mechanisms. Rather than solving the issue with minimal, targeted changes, OH can enter error-prone iterative loops that result in massive file rewrites or reimplementation of the existing core software components. These behaviors likely stem from the limitations of the agentic process in comprehending the full context of the issue and the project. As a result, OH may generate excessive or redundant code that is potentially vulnerable.
@@
In contrast, the OpenHands’ patch bypassed Django’s ORM entirely and rewrote the validation logic using raw SQL:
@@
Vulnerability Types and Overlaps: Among the agentic frameworks, OH, responsible for 44 vulnerabilities, exhibited a similar vulnerability pattern to that of Llama 3.3: command injection (CWE-78, n=10), eval injection (CWE-95, n=8), and insecure deserialization (CWE-502, n=6) were all among the common CWEs in OH patches. This alignment suggests that even when LLMs are embedded within frameworks, some core security risks and patterns persist.
@@
Moreover, we find that vulnerabilities introduced by the agentic frameworks are entirely different from those written by the developers. These results echo the same observations in Section 3.1. While developers, Llama, and agentic frameworks all can add vulnerabilities to the codebase, the vulnerabilities produced by the LLMs or agentic frameworks are not simply repeating the developers’ mistakes.
RQ2 Summary: While the agentic frameworks were evaluated on a smaller subset of data (test-set), all three still introduced a number of vulnerabilities. OH, in particular, produced an order of magnitude more vulnerabilities than ACR or HC. These results indicate that, despite improvements in autonomy and reasoning, current agentic workflows remain vulnerable to security issues, especially when given full control over the codebase.
3.3. RQ3: What code, issue, or project characteristics are associated with the generation of vulnerable code by LLMs when resolving issues?
3.3.1. Code-Level Factors
We analyze the code-level characteristics of the files associated with vulnerability-inducing LLM patches and compare them to those from all LLM-modified or generated files. Table 3 summarizes these comparisons for Llama-generated patches (RQ1). We note that the dataset for RQ1 is significantly larger than RQ2. As a result, due to the limited number of vulnerable examples in ACR and HC, we report statistical comparisons only for OH. However, we include the raw values for all three frameworks in Table 4.